I drag a mass m=21 kg in a straight line, along a horizontal surface, a distance D= 23m .I drag it at constant speed v=0.90m/s in a straight line using a horizontal force.the coefficient of static friction is 1.2 and coefficient of kinetic friction is 1.1.How much work do I do in joules?
1
Expert's answer
2020-05-18T10:22:00-0400

As the object is in motion hence kinetic friction will take place

kinetic friction = "\mu_k\times N=1.1\times21\times10" ="231 N"

lets apply work energy theorem

"\sum W=\Delta K"

"W_N+W_f+W_{mg}+W_F=\Delta K"

"0-(231\times23)+0+W_F=\frac12\times21\times0.9^2"

"W_F=5321.5 J"